OpenStudy (anonymous):

Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers. I'm not quite sure how to do this. Please help me do it. :)

OpenStudy (anonymous):

x = x has all real numbers as its solution ax = ax (multiply both sides by any real number) ax + b = ax + b (add, or subract, the same real number to both sides For example: 2x + 3 = 2x + 3 -3 -3 2x = 2x --- --- 2 2 x = x If you complicate it more, then it will be more than two steps.

OpenStudy (anonymous):

Ohh so say, 4x + 2 = 4x + 2 would be a 2 step equation with all real numbers?

OpenStudy (anonymous):

@kathy0514

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, Thanks! :)

OpenStudy (anonymous):

So to solve it, it should look like?: @kathy0514 \[4x + 2(-2) = 4x + 2(-2)\]\[4x = 4x\]

OpenStudy (anonymous):

THe parentheses is not for multipulcation

OpenStudy (anonymous):

yes